OpenStudy (konradzuse):

use logarithmic differentiation to find dy/dx y = (5x-3)/(3x+2)^3

OpenStudy (konradzuse):

Apparently from my last question like this I did normal differentiation...

OpenStudy (anonymous):

not sure what logarithmic differentiation is ... it's been a while lol let me check

OpenStudy (anonymous):

take the ln of boht sides

OpenStudy (precal):

why not just use the quotient rule?

OpenStudy (anonymous):

since your doing logarithmic differentiation, yes, @Outkast3r09 , is correct, you'll need to take the ln of both sides...

OpenStudy (anonymous):

the question just asks to do it that way for some reason but whatever you should get \[ln(y)=ln(5x-3)-ln((3x+2)^3)\] used the ln(a/b)=ln(a)-ln(b)

OpenStudy (anonymous):

differentiate both side in respects to x

OpenStudy (precal):



OpenStudy (konradzuse):

so ln(y) = 5/(5x-3) - (3(3x+2)^2)/(3x+2)^3) *3?

OpenStudy (konradzuse):

?

OpenStudy (anonymous):

gotta do both sides the left side is once again chain rule where u =ln(y) \[\frac{du}{dx}=\frac{du}{dy}*\frac{dy}{dx}\]

OpenStudy (anonymous):

the right side is correct though

OpenStudy (konradzuse):

so 1/y * null?

OpenStudy (konradzuse):

since y' of y would be 1 actually...

OpenStudy (konradzuse):

not null

OpenStudy (precal):